# Field transformations in functional integral, effective action and functional flow equations

### v2

@@ Introduction
In the presence of fluctuations the relevant field equations are the ones obtained by variation of the quantum effective action or free energy with respect to the fields. They are exact identities for a given form of the effective action and contain the relevant physical information, in contrast to the field equations derived by variation of the classical action. Field transformations of the effective action, that we call here “macroscopic field transformations”, lead to identical field equations, simply related by a variable transformations in differential equations. The quantum effective action or free energy contains much more information than only the field equations. The second functional derivative is related to the connected correlation function or inverse propagator, and interaction vertices follow from higher functional derivatives. After a field transformation the higher functional derivatives of the effective action may no longer be directly related to microscopic correlation functions. They still play an important role for macroscopic observations which are often not sensitive to the microscopic details or microscopic correlation functions. A well known example is the universality of critical phenomena where vastly different microscopic theories can lead to the same free energy or effective action. The second functional derivative of the effective action contains the information about the correlation length independently of the microscopic model. We name the functional derivatives of the effective action “macroscopic correlation functions” and discuss how their relation to microscopic correlation functions is affected by field transformations. This paper aims for an overview of the impact of field transformations on different levels. In the view of recent discussions we put particular emphasis on the technical simplifications that can be realized by field transformations in functional flow equations.
@@
Field transformations can help to focus on important physical features and to discard redundant information. For critical phenomena in statistical physics their use has a long tradition Wegner (1974). Various versions of field transformations have been formulated for modern functional flow equations Wetterich (1996); Gies and Wetterich (2002, 2004); Pawlowski (2007); Floerchinger and Wetterich (2009). They are used for making important “composite or collective degrees of freedom” of a model directly accessible or for technical simplifications. Motivated by the elimination of redundant couplings for the counting of relevant parameters at a given fixed point of renormalization group equations Wegner (1974), the proposal of the “essential renormalization group” Baldazzi et al. (2022); Baldazzi and Falls (2021) has argued that couplings which can be removed by field transformations are not observable and need not to be considered for the search of possible scaling solutions. The essential renormalization group entails important technical simplifications for approximate solutions of functional flow equations Knorr (2023a); Knorr et al. (2022a); Laporte et al. (2023); Ihssen and Pawlowski (2023a); Buccio and Percacci (2022); Knorr et al. (2022b); de Brito et al. (2023); Saueressig (2023); Ihssen and Pawlowski (2023b); Kawai and Ohta (2023); Knorr (2023b); Baldazzi et al. (2023). A critical discussion of the merits and shortcomings of this approach seems timely.
@@
Our second example points to the opposite direction. The inverse propagator of electrons in a solid is of the form $P_{F}(q)=\varepsilon(q)-\mu+i(2n+1)\pi T$, where $\mu$ is the chemical potential, $T$ the temperature and the integer $n$ determines the Matsubara frequency. The Fermi surface is determined by $\varepsilon(q_{F})=\mu$. It is located at non-zero Fermi momenta $q_{F}$ even for the simplest case $\varepsilon(q)=q^{2}/(2m)$. In solids $\varepsilon(q)$ can have a rather complex form which fixes the detailed location of the Fermi surface and many other directly observable features. For $T\neq 0$ a regular linear field transformation transforms $P_{F}(q)\to\tilde{P}_{F}(q)=f(q)P_{F}(q)$. We can bring $\tilde{P}_{F}(q)$ to an arbitrarily chosen form without zeros. After the field transformation $\tilde{P}_{F}(q)$ does no longer contain the information about the Fermi surface if we take some standard form as $\tilde{P}_{F}=q^{2}/(2m)+i(2n+1)\pi T$ for a gas of free fermions or bosons at zero density. The corresponding complex function $f(q)$ has removed the chemical potential and all structures in the dispersion relation $\varepsilon(q)$. The details of the dispersion relation involves couplings that parameterize the electron propagator in a solid. For the two-dimensional Hubbard model with next-neighbor and next-to-nearest neighbor hopping parameters $t$ and $t^{\prime}$ one has
@@
One of the couplings that is removed by this field transformation is the ratio $t^{\prime}/t$.
The conclusion that couplings used to parameterize the momentum-dependence of the fermion propagator are unobservable is not justified. The physical information contained in these couplings is moved by the field transformation to the vertices. Only certain products of vertices with powers of $P_{F}^{-1/2}(q)$ are invariant under the momentum-dependent field transformation. If one would use a truncation which neglects the detailed momentum-dependence of vertices, key physical information would even be lost by a field transformation which brings the electron propagator to some standard form.
@@
The possibility of field transformations has also important conceptual implications. This is most striking for gravity. A non-linear transformation of the metric field changes the geometry. For example, an expanding universe can be transformed into a shrinking universe Wetterich (2013a); Bars et al. (2013); Wetterich (2014a, b); Li (2014); Hossain et al. (2015); Domènech and Sasaki (2015); Ijjas and Steinhardt (2015); Cai and Piao (2016); Domènech and Sasaki (2016); Kamenshchik et al. (2016); Graef et al. (2017); Amendola et al. (2018); Shimon (2019); Luan and Qiu (2020); Casadio et al. (2021); Koivisto and Zheng (2021); Rubakov and Wetterich (2022); Mukherjee and Sahota (2023); Lombriser (2023). Field relativity Wetterich (2013a) states that macroscopic field transformations of the quantum effective action do not change the observables. Cosmologies with different geometries related by such transformations are strictly equivalent. The present paper confirms this point of view. Field relativity also establishes the equivalence of many models of modified gravity with coupled quintessence Wetterich (2014c).
As a simple example illustrating various aspects of field transformations we focus in this paper on a quantum field theory for a single scalar, with a classical action involving only a term with two derivatives, multiplied by a field-dependent wave function renormalization or “kinetial”. On the one hand a non-linear field transformation can transform away the field-dependence of the kinetial such that the transformed action looks like the one for a free massless scalar field. On the other hand the field-dependence of the kinetial leads to terms in the action involving more than two powers of the field, suggesting interactions. We will relate these various aspects and find that for generic kinetials this model indeed describes an interacting scalar field.
@@
This describes a free massless composite field ${\varphi}\sim\ln\chi$. This ultraviolet fixed point differs from the (trivial) infrared fixed point for which the kinetial is field-independent. The ultraviolet fixed point is characterized by the quantum scale symmetry of a constant multiplicative rescaling $\chi\to\beta\chi$ with fixed metric. In contrast, the infrared fixed point exhibits a different form of quantum scale symmetry for which the metric (or coordinates) are rescaled together with $\chi$. It also is invariant under constant shifts $\chi\to\chi+\delta$. Possible crossover trajectories between the two fixed points are characterized by the presence of interactions. If the ultraviolet fixed point (2) is confirmed this would define a possible ultraviolet completion of an interacting scalar field theory in four dimensions, avoiding the “triviality problem”.
In sect. II we begin with field-transformations for the integration variables of the functional integral defining the quantum field theory. The Jacobian of this transformation modifies the action beyond a simple replacement of the field variables. We demonstrate how non-linear field transformations change the saddle point approximation and perturbation theory. Sect. III turns to field transformations for the quantum effective action for which all effects of quantum fluctuations are included. We distinguish between macroscopic field transformations which only modify the expression of a given effective action in terms of field-variables, and microscopic field transformations which modify in addition the sources used for the definition of the effective action.
In sect. IV we proceed to the field transformations in functional flow equations, focusing first on the macroscopic field transformations Wetterich (1996); Gies and Wetterich (2002). This is contrasted in sect. V with the microscopic field transformations Pawlowski (2007); Baldazzi et al. (2022); Baldazzi and Falls (2021); Ihssen and Pawlowski (2023b) which are the basis of the proposed “essential renormalization group”. In sect. VI we employ a formulation with both fundamental and composite fields in order to clarify further aspects of the microscopic field transformations, in particular the relation between the classical action and the initial values of the flow equations. Sect. VII concentrates on scaling solutions of the flow equations whose existence can make a model ultraviolet complete. Our conclusions are presented in sect. VIII. Two appendices provide a discussion of the model with a field-dependent kinetial without the use of field transformations. This can be used for comparison with treatments involving field transformations. These computations yield indications for the exitence of the ultraviolet fixed point (2).
@@
The issue of a change of field-variables is rather different on the level of the (quantum) effective action $\Gamma$. The computation of $\Gamma$ performs already the functional integral. Once the effective action is computed or its form assumed, no further integration over fields is needed and the Jacobian for integration variables plays no longer a role. The quantum field equation and the macroscopic correlation functions are obtained by taking functional derivatives of the effective action. On this level one can perform arbitrary transformations of field variables. Physical observables are often expressed in terms of the macroscopic correlation functions, which in turn are given by functional derivatives of $\Gamma$. The expectation values of observables remain the same if the field transformations are inserted in the functional derivatives of the effective action. Field equations are differential equations. They can be solved for an arbitrary choice of field-variables. In short, the physics remains invariant under an arbitrary change of macroscopic field-variables on which a given effective action depends. We call the transformations of the variables in the effective action “macroscopic field transformations”. What is affected by the macroscopic field transformations is the relation between the macroscopic correlation functions and the microscopic correlation functions defined by expectation values of powers of microscopic fields. We discuss this issue in detail.
@@
This expression involves higher correlation functions for $\chi^{\prime}$. The expressions (51) and (III) differ even if one uses the same functions for $N_{m}$ and $N$. Their explicit relation is not obvious. A further difference arises for $\partial\Gamma/\partial\chi(y)\neq 0$.
@@
The functional derivatives of $\Gamma_{\chi}[{\varphi}]$ with respect to ${\varphi}$ do not generate the correlation functions for the microscopic field ${\varphi}^{\prime}$ which obtains from $\chi^{\prime}$ by an analogous field transformation. They rather multiply the correlation functions for $\chi^{\prime}$ with suitable functions of the expectation value $\chi$. In this sense $\chi$ is singled out as the “canonical field” for the effective action $\Gamma_{\chi}$. Fields ${\varphi}[\chi]$ obtained by a non-linear field transformation are no longer canonical as far as the relation of functional derivatives of $\Gamma_{\chi}[{\varphi}]$ to correlation functions of microscopic fields are concerned. They may be called “composite fields”.
@@
We may define “macroscopic correlation functions” for ${\varphi}$ by suitable functional derivatives of $\Gamma_{\chi}$ with respect to ${\varphi}(x)$. For example, the macroscopic connected two-point function for ${\varphi}$ is given by the inverse of $\Gamma_{({\varphi})}^{(2)}$ in eq. (50). In contrast, the “microscopic correlation functions” are given by suitable expectation values of powers of the microscopic field ${\varphi}^{\prime}(x)$. For the canonical field $\chi^{\prime}$ the macroscopic and microscopic correlation functions coincide. For the non-canonical field ${\varphi}^{\prime}$ we have seen that in general the macroscopic and microscopic correlation functions differ.
For many problems the knowledge of the microscopic correlation function plays actually no role. Only the macroscopic correlation functions matter, and therefore only suitable macroscopic variable transformations in the effective action $\Gamma_{\chi}$ are involved. As an example, we may consider the field equation in the presence of a source,
@@
The role of the macroscopic correlation functions is particularly apparent for critical systems. Many, sometimes vastly different, microscopic models can lead to the same universal critical behavior which is encoded in the form of the free energy of effective action $\Gamma$. Important quantities as the correlation length can be extracted directly from $\Gamma$ for an arbitrary choice of fields, reflecting properties of the second functional derivative $\Gamma^{(2)}$. The microscopic correlation functions play no role here – they can be actually very different for the different microscopic models. Only the macroscopic fields play a role for the universal critical behavior. In terms of the microscopic fields of a given microscopic model they are often composite fields.
@@
Different versions of the quantum effective action correspond to different choices of source terms. They are related by a type of microscopic frame transformation which also transforms the source term. These different versions are strictly equivalent as far as solutions of the quantum field equations in the absence of sources are concerned. They often remain (approximately) equivalent for the macroscopic correlation functions as well.
@@
At this point it may be useful to highlight an important difference between macroscopic and microscopic field transformations. The macroscopic field transformations do not change the effective average action. It only expresses a given functional in terms of different field-variables. In this way one can explicitly follow the action of the transformation once a given approximation to $\Gamma_{k}$ is given. For example, a transformation ${\varphi}(x)=f^{-1}_{k}\big{(}\chi(x)\big{)}$ transforms $U(\chi)\to U({\varphi})=U\big{(}f_{k}({\varphi})\big{)}$. A regular transformation $f_{k}$ can transform an arbitrary convex potential $U(\chi)$ to $U({\varphi})=m^{2}{\varphi}^{2}+U_{0}$ for a suitable range of fields. This holds for arbitrarily small $m^{2}$. A potential with spontaneous symmetry breaking of the discrete symmetry $\chi\to-\chi$ cannot be transformed to this form, however. Monotonically decreasing potentials can be brought to the form $U=M^{4}\exp(-{\varphi}/M)$ for a suitable range of fields. Transforming the potential to a standard form shifts information to the kinetial. Standard forms of the potential have proven to be useful for inflationary cosmology Wetterich (2013b); Kallosh et al. (2014); Wetterich (2015c); Galante et al. (2015) or dynamical dark energy Hebecker and Wetterich (2001); Wetterich (2003).
@@
Conclusions

### v3

@@ Introduction
In the presence of fluctuations the relevant field equations are the ones obtained by variation of the quantum effective action or free energy with respect to the fields. They are exact identities for a given form of the effective action and contain the relevant physical information, in contrast to the field equations derived by variation of the classical action. Field transformations of the effective action, that we call here “macroscopic field transformations”, lead to identical field equations, simply related by a variable transformations in differential equations. The quantum effective action or free energy contains much more information than only the field equations. The second functional derivative is related to the connected correlation function or inverse propagator, and interaction vertices follow from higher functional derivatives. After a field transformation the higher functional derivatives of the effective action may no longer be directly related to microscopic correlation functions. They still play an important role for macroscopic observations which are often not sensitive to the microscopic details or microscopic correlation functions. A well known example is the universality of critical phenomena where vastly different microscopic theories can lead to the same free energy or effective action. The second functional derivative of the effective action contains the information about the correlation length independently of the microscopic model. We name the functional derivatives of the effective action “macroscopic correlation functions” and discuss how their relation to microscopic correlation functions is affected by field transformations.
This paper aims for an overview of the impact of field transformations on different levels. In the view of recent discussions we put particular emphasis on the technical simplifications that can be realized by field transformations in functional flow equations.
@@
Field transformations can help to focus on important physical features and to discard redundant information. For critical phenomena in statistical physics their use has a long tradition Wegner (1974). Various versions of field transformations have been formulated for modern functional flow equations Wetterich (1996); Gies and Wetterich (2002, 2004); Pawlowski (2007); Floerchinger and Wetterich (2009). They are used for making important “composite or collective degrees of freedom” of a model directly accessible or for technical simplifications. Motivated by the elimination of redundant couplings for the counting of relevant parameters at a given fixed point of renormalization group equations Wegner (1974), the proposal of the “essential renormalization group” Baldazzi et al. (2022); Baldazzi and Falls (2021) has argued that couplings which can be removed by field transformations are not observable and need not to be considered for the search of possible scaling solutions. We question this statement. The essential renormalization group entails important technical simplifications for approximate solutions of functional flow equations Knorr (2023a); Knorr et al. (2022a); Laporte et al. (2023); Ihssen and Pawlowski (2023a); Buccio and Percacci (2022); Knorr et al. (2022b); de Brito et al. (2023); Saueressig (2023); Ihssen and Pawlowski (2023b); Kawai and Ohta (2023); Knorr (2023b); Baldazzi et al. (2023). A critical discussion of the merits and shortcomings of this approach seems timely.
@@
Our second example points to the opposite direction. The inverse propagator of electrons in a solid is of the form $P_{F}(q)=\varepsilon(q)-\mu+i(2n+1)\pi T$, where $\mu$ is the chemical potential, $T$ the temperature and the integer $n$ determines the Matsubara frequency. The Fermi surface is determined by $\varepsilon(q_{F})=\mu$. It is located at non-zero Fermi momenta $q_{F}$ even for the simplest case $\varepsilon(q)=q^{2}/(2m)$. In solids $\varepsilon(q)$ can have a rather complex form which fixes the detailed location of the Fermi surface and many other directly observable features. For $T\neq 0$ a regular linear field transformation transforms $P_{F}(q)\to\tilde{P}_{F}(q)=f(q)P_{F}(q)$. We can bring $\tilde{P}_{F}(q)$ to an arbitrarily chosen form without zeros. After the field transformation $\tilde{P}_{F}(q)$ does no longer contain the information about the Fermi surface if we take some standard form as $\tilde{P}_{F}=q^{2}/(2m)+i(2n+1)\pi T$ for a gas of free fermions or bosons at zero density. The corresponding complex function $f(q)$ has removed the chemical potential and all structures in the dispersion relation $\varepsilon(q)$.
The multiplication of the fermion fields by $f^{-1/2}(q)$ is linear, but highly non-local. Powers of $f^{1/2}(q)$ multiply the vertices once expressed in terms of the rescaled fields. This includes the source term $\int_{q}j(q)\psi(q)=\int_{q}\tilde{j}(q)\tilde{\psi}(q)$, $\tilde{j}(q)=f^{1/2}(q)j(q)$. If the source term is local in the original fermion field $\psi(q)$, it will be non-local for the transformed field $\tilde{\psi}(q)$. The source term is not directly visible in the effective action. Deriving field equations for non-zero sources like the non-local transformation transformation of the source becomes important.
The details of the dispersion relation involves couplings that parameterize the electron propagator in a solid. For the two-dimensional Hubbard model with next-neighbor and next-to-nearest neighbor hopping parameters $t$ and $t^{\prime}$ one has
@@
One of the couplings that is removed by the above field transformation is the ratio $t^{\prime}/t$. The conclusion that couplings used to parameterize the momentum-dependence of the fermion propagator are unobservable is not justified even for vanishing sources. The physical information contained in these couplings is moved by the field transformation to the vertices. Only certain products of vertices with powers of $P_{F}^{-1/2}(q)$ are invariant under the momentum-dependent field transformation. If one would use a truncation which neglects the detailed momentum-dependence of vertices, key physical information would even be lost by a field transformation which brings the electron propagator to some standard form.
@@
The possibility of field transformations has also important conceptual implications. This is most striking for gravity. A non-linear transformation of the metric field changes the geometry. For example, an expanding universe can be transformed into a shrinking universe Wetterich (2013a); Bars et al. (2013); Wetterich (2014a, b); Li (2014); Hossain et al. (2015); Domènech and Sasaki (2015); Ijjas and Steinhardt (2015); Cai and Piao (2016); Domènech and Sasaki (2016); Kamenshchik et al. (2016); Graef et al. (2017); Amendola et al. (2018); Shimon (2019); Luan and Qiu (2020); Casadio et al. (2021); Koivisto and Zheng (2021); Rubakov and Wetterich (2022); Mukherjee and Sahota (2023); Lombriser (2023). Field relativity Wetterich (2013a) states that macroscopic field transformations of the quantum effective action do not change the observables. Cosmologies with different geometries related by such transformations are strictly equivalent. Of course, this necessitates the simultaneous transformation of the energy momentum tensor as a source term Wetterich (1988). The present paper confirms this point of view. Field relativity also establishes the equivalence of many models of modified gravity with coupled quintessence Wetterich (2014c).
As a simple example illustrating various aspects of field transformations we focus in this paper on a quantum field theory for a single scalar field, with a classical action involving only a term with two derivatives, multiplied by a field-dependent wave function renormalization or “kinetial”. On the one hand a non-linear field transformation can transform away the field-dependence of the kinetial such that the transformed action looks like the one for a free massless scalar field. On the other hand the field-dependence of the kinetial leads to terms in the action involving more than two powers of the field, suggesting interactions. We will relate these various aspects and find that for generic kinetials this model indeed describes an interacting scalar field.
@@
This describes a free massless composite field ${\varphi}\sim\ln\chi$. This ultraviolet fixed point differs from the (trivial) infrared fixed point for which the kinetial is field-independent. The ultraviolet fixed point is characterized by the quantum scale symmetry of a constant multiplicative rescaling $\chi\to\beta\chi$ with fixed metric. In contrast, the infrared fixed point exhibits a different form of quantum scale symmetry for which the metric (or coordinates) are rescaled together with $\chi$. The effective action for the infrared fixed point is invariant under constant shifts $\chi\to\chi+\delta$. Possible crossover trajectories between the two fixed points are characterized by the presence of interactions. If the ultraviolet fixed point (2) is confirmed this would define a possible ultraviolet completion of an interacting scalar field theory in four dimensions, avoiding the “triviality problem”.
In sect. II we begin with field-transformations for the integration variables of the functional integral defining the quantum field theory. The Jacobian of this transformation modifies the action beyond a simple replacement of the field variables. We discuss how field transformations affect the relation between macroscopic and microscopic correlation functions. On the level of the effective action we demonstrate how non-linear field transformations change the saddle point approximation and perturbation theory. Sect. III turns to field transformations for the quantum effective action for which all effects of quantum fluctuations are included. We distinguish between macroscopic field transformations which only modify the expression of a given effective action in terms of field-variables, and a type of microscopic field transformations which modify in addition the sources used for the definition of the effective action.
In sect. IV we proceed to the field transformations in functional flow equations, focusing first on the macroscopic field transformations Wetterich (1996); Gies and Wetterich (2002). This is contrasted in sect. V with the microscopic field transformations Pawlowski (2007); Baldazzi et al. (2022); Baldazzi and Falls (2021); Ihssen and Pawlowski (2023b) which are the basis of the proposed “essential renormalization group”. In sect. VI we employ a formulation with both fundamental and composite fields in order to clarify further aspects of the microscopic field transformations, in particular the relation between the classical action and the initial values of the flow equations. Sect. VII concentrates on scaling solutions of the flow equations whose existence can make a model ultraviolet complete. A conclusding discussion is presented in sect. VIII. Two appendices provide a discussion of the model with a field-dependent kinetial without the use of field transformations. This can be used for comparison with treatments involving field transformations. These computations yield indications for the exitence of the ultraviolet fixed point (2).
@@
The quantum effective action $\Gamma$ is a very strong concept in quantum field theory. Since it incorporates all effects of quantum fluctuations. The analogue in equilibrium statistical physics is a field-dependent Gibbs free energy which includes all effects of thermal fluctuations. We will use the common wording “effective action” for both, as well as for similar objects for which fluctuation effects of a functional integral or probability distribution are fully included. The macroscopic field equations obtain by setting the first functional derivative of the effective action equal to a source. They are exact identities without any further corrections. For macroscopic observations these are the only relevant “classical field equations” in the context of fluctuating microphysics. For gravity the source is the energy momentum tensor, for electrodynamics in medium an external charge distribution, or for superconductivity an external magnetic field.
For a given solution of the field equation one may investigate a small perturbation of the source. The response of the solution to this perturbation in found by linearizing the field equation around the given solution for the unperturbed source. This involves the second functional derivative $\Gamma^{(2)}$ of the effective action. The inverse of $\Gamma^{(2)}$ defines the propagator as well known for the general setting of Green’s functions. In our context we call this propagator a “macroscopic correlation function”. Higher functional derivatives of $\Gamma$ define vertices and the associated coupling constants for effective interactions.
Macroscopic observables are expressed in terms of functional derivatives of the effective action. This relation is typically formulated for a given choice of fields. It may be rather complex, as the extraction of the scattering matrix from the effective action. For our purposes it is sufficient that all relevant macroscopic observables can be computed in principle as functions of the functional derivatives of the effective action. The effective action contains then the full information about all relevant physical observables.
The effective action will not be known exactly except for special cases as free field theories. Nevertheless, powerful approximations (“truncations”) exist based on symmetries and expansions in the number of derivatives, small couplings etc. Quite often one simply assumes a valid approximation for the effective action. This is the case for Landau type theories for superconductors or superfluids. From this ansatz for the effective action many physical observables are computed. For example, the macroscopic two-point function or propagator is extracted from the second functional derivative of the effective action. It contains the correlation lengths for superconductors or the Fermi surface or correlated electrons in solids. This computation of the correlation length requires a local coupling of the field to the source, which is usually assumed tacitly.
In a quantum field theory context general relativity is based on the assumption that the effective action for a macroscopic metric field takes the form of the Einstein-Hilbert action with a cosmological constant. This is dictated by diffeomorphism symmetry and an expansion in second order in derivatives. We emphasize that this assumption concerns the effective action. The underlying microscopic degrees of freedom of a theory of quantum gravity are not known. For given microscopic degrees of freedom we do not know their precise relation to the macroscopic metric. Also the classical action for some assumed microscopic degrees of freedom of quantum gravity is not known. Nevertheless, a huge amount of observables is derived from the assumed form of the effective action.
More generally, on the level of the effective action the direct contact to microphysics may be lost. For example, the Kosterlitz-Thouless phase transition occurs for many very different two-dimensional systems with continuous $\text{SO}(2)$-symmetry. The universal critical behavior is encoded in an ansatz for the effective action containing only a few functions defining the universal equation of state. From this ansatz one can compute the macroscopic correlation function and many other quantities. The precise relation to the microscopic correlation functions for the different models realizing the Kosterlitz-Thouless transition may be rather involved and is often not known in practice. Nevertheless, the macroscopic correlation functions describe a good deal of the relevant macroscopic observations.
Macroscopic field transformations are variable transformations of the field-arguments of the effective action. For a given effective action they can be performed without the knowledge of the underlying microscopic action. Such macroscopic field transformations affect, however, the relation between macroscopic and microscopic correlation functions. If one is interested in the precise connection between macrophysics and microphysics one has to keep track of this effect. This also concerns the issue of field transformations in flow equations that we will discuss in the following sections.
In order to shed light on the precise relation between macrophysics and microphysics we investigate in this section the computation of the effective action from the classical action of a scalar quantum field theory. In this context we can address the question what are the relation between macroscopic and microscopic correlation functions and how they behave under field transformations. For a choice of “canonical fields” the relation between macroscopic and microscopic correlation function follows directly from the definition of the effective action. They are identical. Performing non-linear microscopic or macroscopic field transformations this direct connection between macroscopic and microscopic correlation functions will be lost. Often we do not know the precise microscopic composite field whose expectation value yields the macroscopic field. Nor do we know which precise choice of macroscopic field corresponds to a canonical field. Our discussion highlights the complexity of the precise relation between macrophysics and microphysics.
@@
The issue of the macroscopic change of field variables for the effective action $\Gamma$ differs conceptually from the microscopic field transformation on the level of the functional integral discussed in the previous section. The computation of $\Gamma$ performs already the functional integral. Once the effective action is computed or its form assumed, no further integration over fields is needed and the Jacobian for integration variables plays no longer a role. The quantum field equation and the macroscopic correlation functions are obtained by taking functional derivatives of the effective action. On this level one can perform arbitrary transformations of field variables. They translate directly to the functional derivatives of the effective action by the standard rules for differentiation. By our assumption physical observables are expressed as functions of the functional derivatives of the effective action or the macroscopic correlation functions. The expectation values of observables remain the same if the field transformations are inserted in the functional derivatives of the effective action. Of course, the relation between observables and macroscopic correlation functions change their functional form by this insertion.
Field equations are differential equations. They can be solved for an arbitrary choice of field-variables. Or course, for non-vanishing sources on has to take the transformation of the source term into account. It transforms as $\partial\Gamma/\partial{\varphi}$. For a transformation from ${\varphi}$ to $\tilde{{\varphi}}$, defined by a function ${\varphi}(\tilde{{\varphi}})$, the transformed source $\tilde{j}$ for $\tilde{{\varphi}}$ obeys $\tilde{j}=(\partial{\varphi}/\partial\tilde{{\varphi}})j$, with $j$ the source for ${\varphi}$.
In short, the physics remains invariant under an arbitrary change of macroscopic field-variables on which a given effective action depends. We call the transformations of the variables in the effective action “macroscopic field transformations”. What is affected by the macroscopic field transformations is the relation between the macroscopic correlation functions and the microscopic correlation functions defined by expectation values of powers of microscopic fields. We discuss this issue in detail.
@@
For the “canonical” field $\chi$ the macroscopic and the microscopic correlation functions coincide. This is visible, for example, in the relation (45) for which $(\Gamma^{(2)})^{-1}$ is the macroscopic connected two-point function and $G$ the microscopic connected two-point function. For non-linear field transformations this direct relation between macroscopic and microscopic correlations is lost. In general, there is no longer any simple relation between the macroscopic correlations for the transformed field and the microscopic correlations for some field obtained by a microscopic field transformation. The direct relation between macroscopic and microscopic correlations is maintained only for linear field transformations.
@@
For non-linear field transformations this expression involves higher correlation functions for $\chi^{\prime}$. The expressions (51) and (III) differ even if one uses the same functions for $N_{m}$ and $N$. Their explicit relation is not obvious. A further difference arises for $\partial\Gamma/\partial\chi(y)\neq 0$. For non-linear field transformations with field dependent $N_{m}$ and $N$ the relation between macroscopic and microscopic correlations is rather complex. The relation (III) for the microscopic correlation function for ${\varphi}^{\prime}$ contains higher correlations for $\chi^{\prime}$. They can be expressed in terms of functional derivatives of $\Gamma$ with respect to the canonical field $\chi$. In turn, these macroscopic correlations for $\chi$ can be expressed in terms of macroscopic correlations for ${\varphi}$. In the other direction, the macroscopic correlation function $(\Gamma_{({\varphi})}^{(2)})^{-1}$ involves the microscopic two-point function for $\chi^{\prime}$ and functions of expectation values $\langle\chi^{\prime}\rangle=\chi$. We can express $\chi$ in terms of ${\varphi}$ by the field transformation. The complication of a relation to microscopic correlations for ${\varphi}^{\prime}$ arises for the complicated relation between correlations for ${\varphi}^{\prime}$ and $\chi^{\prime}$, as in eq. (III). In summary, if one wants to keep precise track of microscopic correlations the use of non-linear field transformations is perhaps not a very good idea.
@@
The functional derivatives of $\Gamma_{\chi}[{\varphi}]$ with respect to ${\varphi}$ do not generate the correlation functions for the microscopic field ${\varphi}^{\prime}$ which obtains from $\chi^{\prime}$ by an analogous field transformation. They rather multiply the correlation functions for $\chi^{\prime}$ with suitable functions of the expectation value $\chi$. In this sense $\chi$ is singled out as the “canonical field” for the effective action $\Gamma_{\chi}$. Fields ${\varphi}[\chi]$ obtained by a non-linear field transformation are no longer canonical as far as the relation of functional derivatives of $\Gamma_{\chi}[{\varphi}]$ to correlation functions of microscopic fields are concerned.
@@
We define “macroscopic correlation functions” for ${\varphi}$ by suitable functional derivatives of $\Gamma_{\chi}$ with respect to ${\varphi}(x)$. For example, the macroscopic connected two-point function for ${\varphi}$ is given by the inverse of $\Gamma_{({\varphi})}^{(2)}$ in eq. (50). In contrast, the “microscopic correlation functions” are given by suitable expectation values of powers of the microscopic field ${\varphi}^{\prime}(x)$. For the canonical field $\chi^{\prime}$ the macroscopic and microscopic correlation functions coincide. For the non-canonical field ${\varphi}$ or ${\varphi}^{\prime}$ we have seen that in general the macroscopic and microscopic correlation functions differ even if we choose the same function $N_{m}$ and $N$.
As we have argued at the beginning of this section for many problems the knowledge of the microscopic correlation function plays actually no role. Only the macroscopic correlation functions matter, and therefore only suitable macroscopic variable transformations in the effective action $\Gamma_{\chi}$ are involved. As an example, we may consider the field equation in the presence of a source,
@@
The role of the macroscopic correlation functions is particularly apparent for critical systems. Many, sometimes vastly different, microscopic models can lead to the same universal critical behavior which is encoded in the form of the free energy of effective action $\Gamma$. Important quantities as the correlation length can be extracted directly from $\Gamma$ for an many choices of fields, reflecting properties of the second functional derivative $\Gamma^{(2)}$. This only needs locality of the coupling of fields to sources. The latter is maintained by local non-linear field transformations. The microscopic correlation functions play no role here – they can be actually very different for the different microscopic models. Only the macroscopic fields play a role for the universal critical behavior. In terms of the microscopic fields of a given microscopic model they are often rather complex composite fields.
@@
Different versions of the quantum effective action correspond to different choices of source terms. They are related by a type of microscopic frame transformation which also transforms the source term. These different versions are strictly equivalent as far as solutions of the quantum field equations in the absence of sources are concerned. They often remain (approximately) equivalent for the macroscopic correlation functions as well. Nevertheless, on the conceptional level they differ. The microscopic field transformations change the effective action, while the fields remain canonical. The macroscopic field transformations leave the effective action invariant, while a canonical field does not remain canonical.
@@
At this point it may be useful to highlight an important difference between macroscopic and microscopic field transformations. The macroscopic field transformations do not change the effective average action. It only expresses a given functional in terms of different field-variables. In this way one can explicitly follow the action of the transformation once a given approximation to $\Gamma_{k}$ is given. For example, a transformation ${\varphi}(x)=f^{-1}_{k}\big{(}\chi(x)\big{)}$ transforms $U(\chi)\to U({\varphi})=U\big{(}f_{k}({\varphi})\big{)}$. A regular transformation $f_{k}$ can transform an arbitrary convex potential $U(\chi)$ to $U({\varphi})=m^{2}{\varphi}^{2}+U_{0}$ for a suitable range of fields. This holds for arbitrarily small $m^{2}$. A potential with spontaneous symmetry breaking of the discrete symmetry $\chi\to-\chi$ cannot be transformed to this form, however. Monotonically decreasing potentials can be brought to the form $U=M^{4}\exp(-{\varphi}/M)$ for a suitable range of fields. Transforming the potential to a standard form shifts information to the kinetial. Standard forms of the potential have proven to be useful for inflationary cosmology Wetterich (2013b); Kallosh et al. (2014); Wetterich (2015b); Galante et al. (2015) or dynamical dark energy Hebecker and Wetterich (2001); Wetterich (2003).
@@
Discussion
The present paper addresses a systematic discussion of field transformations, starting from variables in the functional integral and proceeding to transformations of field variables in the effective action and for flow equations for associated scale-dependent effective average actions. In order to illustrate the conceptual discussion with a practical case we have focused often on a simple scalar model with a field-dependent kinetic term. For our concluding discussion we focus on field transformations for functional flow equations.